There are twenty dogs in the “Top Dog Show”. Three of the dogs will be chosen as finalists for top honors. There are how many possible ways for a dog to be selected as a finalist?
Answer by onlinepsa(22)

.
.
Number of ways 3 dogs cab be chosen from 20 to be the finalists= 20C3

Number of ways a dog can be chosen from the chosen 3 finalists= 20C3*3C1= 20*19*18/2= 3420.



Thanks,
PRD
https://onlinepsa.wordpress.com/